Jim can fill a pool carrying bucks of watter in 30minutes. Sue can do the job in 45 minutes. Tony can do the same job in 1 hr and 30 minutes. How quickly can all three fill the pool together?
Answer by stanbon(75887)

Jim can fill a pool carrying bucks of water in 30 minutes. Sue can do the job in 45 minutes. Tony can do the same job in 1 hr and 30 minutes. How quickly can all three fill the pool together?
---------------------------------------
Jim rate: 1/30 job/min
Sue rate: 1/45 job/min
Tony rate: 1/90 job/min
Together rate: 1/x job/min
-----------------------------
EQUATION:
1/30 + 1/45 + 1/90 = 1/x
3x + 2x + x = 90
6x = 90
x = 15 (time to fill the pool if they work together)
=================
Cheers,
Stan H.